OpenStudy (anonymous):

can somebody help me integrate (3+x)e^(-x)???i need as soon as possible

OpenStudy (amistre64):

just multiply it out and work it in pieces

OpenStudy (amistre64):

eiterh way youll have to play with integration by parts for a little bit

OpenStudy (amistre64):

say, u = 3+x ; du = dx u-3 = x integrate u e^(u-3) du

OpenStudy (amistre64):

i forgot that negative in your exponent u e^(3-u) du

OpenStudy (amistre64):

that still has a by parts in it, so either way you go your gonna have to get your hands dirty

OpenStudy (anonymous):

so if my answer is [e^{-x} +xe^{-x} + c \] is wrong?

OpenStudy (amistre64):

dunno, i would have to see how you worked it out

OpenStudy (amistre64):

im sure theres a 3 in there someplace thats missing

OpenStudy (amistre64):

i good dbl chk is to take the derivative of what youve got and see if it becomes the stuff you had to begin with

OpenStudy (amistre64):

[e^{-x} +xe^{-x} + c ]' = -e^(-x) - x e^(-x); and that dont equal 3e^(-x) + xe^(-x)

OpenStudy (anonymous):

u know some links that shows the similar example?i need to find out this answer quick

OpenStudy (amistre64):

i forgot to do a product on the middle, but still

OpenStudy (amistre64):

you can find it out real quick by running thru it right now, its simple enough

OpenStudy (anonymous):

i dunno how =(

OpenStudy (amistre64):

how did you get what you posted?

OpenStudy (anonymous):

i use u=3+x,,differentiate got 1,and for the v,i integrate e^(-x),got -e^(-x) then (3+x) multiply with e^(-x)

OpenStudy (amistre64):

e^-x (3+x) -e^-x -1 e^-x -e^-x(3+x-1) = -e^-x (2+x) + c id forget the usub stuff; it doesnt make the integrating simpler for this one

OpenStudy (amistre64):

that there is a table method that follows the integration by parts

OpenStudy (amistre64):

another way to write the answer would be: -2e^-x -x e^-x + c

OpenStudy (amistre64):

im off as well, musta typoed something

OpenStudy (anonymous):

i'm still in confuse right now,,do i have to use certain formua for this kind of integration?

OpenStudy (amistre64):

-e^-x(3+x-1) = -e^-x (2+x) + c ^ there it is, needs to be + -e^-x(3+x+1) = -e^-x (4+x) + c

OpenStudy (amistre64):

its just integration by parts

OpenStudy (amistre64):

i see what you did in yours, you did by parts and i read it as a usub

OpenStudy (anonymous):

ooo i think i should improve more on my math skills like you

OpenStudy (amistre64):

u = 3+x v = -e^-x du = dx dv = e^-x dx \[\int udv=uv-\int vdu\] \[\int (3+x)e^{-x}dx=-(3+x)e^{-x}-\int -e^{-x}dx\] \[\int (3+x)e^{-x}dx=-(3+x)e^{-x}-e^{-x}+c\]

OpenStudy (amistre64):

so, your concept was onteh right track, you just need to hone your skills ;)

OpenStudy (anonymous):

he2,,so that's the final answer?you got it right? =D

OpenStudy (amistre64):

http://www.wolframalpha.com/input/?i=derivative+-4e%5E-x+-x+e%5E-x the wolf likes it

OpenStudy (amistre64):

simplify as wanted

OpenStudy (anonymous):

the wolf needs to help 'chickens' like me more